OpenStudy (anonymous):

3/8 divide by 1/4 = A. 3/32 B. 3/12 C. 1/3 I. 3/2 ~Please show work =)~

OpenStudy (anonymous):

Multiply by the reciprocal of the bottom fraction. So Multiply (3/8) by (4/1)

OpenStudy (anonymous):

3/2 BECAUSE 3/8/1/4=(3/8 )*4=3/2